# Sparse Autoencoders for Hypothesis Generation

### v2

@@ 1 Introduction
Automated approaches to answering these questions have the potential to significantly expand and accelerate scientific discovery (Ludwig & Mullainathan, 2023). Indeed, multiple lines of work—spanning decades of research—have sought to extract text features that can predict a target variable (Blei et al., 2003; Monroe et al., 2009). Recent LLM-based hypothesis generation methods are especially exciting since they operate directly at the level of human-interpretable natural language concepts (Zhou et al., 2024; Batista & Ross, 2024; Ludan et al., 2024; Zhong et al., 2024). Concretely, hypothesis generation methods take a dataset of texts linked to a target variable (e.g., headlines and their engagement level) and hypothesize natural language concepts that predict the target variable (e.g., mentions being surprised or shocked or asks a question to the reader).
@@ 2.1 Hypothesis Generation
Recently, several LLM-based methods aim to automate interpretable concept discovery from text datasets. Grootendorst (2024) proposes using LLMs to assign topics an interpretable label, while Pham et al. (2024) use LLMs for end-to-end topic modeling. Ludan et al. (2024) directly prompt LLMs with a task description to propose interpretable concepts. Sun et al. (2024) and Zhong et al. (2024) also use LLMs to propose concepts, guided by neurons from a supervised model. Zhou et al. (2024) and Batista & Ross (2024) focus specifically on hypothesis generation: LLMs propose natural language descriptions of predictive relationships, which can then be tested for generalization.
@@ 3 Theoretical Framework
Intuitively, the result shows that if $\hat{Z}$ is a high-fidelity predictor of $Z$, then it will also have a similar separation score. A full proof is given in Appendix E. [^2] There are two main observations that enable us to show the result. First, we focus on $|\mathbb{E}[Y|\hat{Z}=1]-\mathbb{E}[Y|Z=1]|,$ and show that if the false negative rate (FNR) is greater than the false discovery rate (FDR), this value is bounded by the FNR; otherwise, it is bounded by the FDR. Second, we show roughly that as long as $\Pr[\hat{Z}=1],\Pr[Z=1]<\Pr[\hat{Z}=0],\Pr[Z=0],$ then $|\mathbb{E}[Y|\hat{Z}=0]-\mathbb{E}[Y|Z=0]|<|\mathbb{E}[Y|\hat{Z}=1]-\mathbb{E}[Y|Z=1]|.$
@@ 4 Methods
In the previous section, we established theoretically that we can generate hypotheses by learning interpretable neurons, identifying which neurons are predictive, and then generating high-fidelity interpretations of the neurons. We now describe how we can accomplish these steps in practice.
We now describe the details of HypotheSAEs. Given a dataset of training examples $\{(x_{i},y_{i})\}_{i\in[N]},$ where $x_{i}$ is the input text and $y_{i}$ is the target variable annotation, HypotheSAEs outputs $H$ natural language concepts that serve as hypotheses. The goal is for these natural language concepts to predict the target variable.
@@
$z_{i}$ $=\operatorname{ReLU}(\operatorname{TopK}(W_{\operatorname{enc}}(e_{i}-b_{\mathrm{pre}})+b_{\operatorname{enc}}))$ (5)
@@
Empirically, the dimensions of $z_{i}$ are often highly interpretable, corresponding to human concepts. $M$ sets the total number of concepts learned across the entire dataset, and $k$ sets the number of concepts which can be used to reconstruct each instance. The output of this first step is an $N\times M$ activation matrix, $Z_{\text{SAE}}$.
@@
In this step, we generate high-fidelity interpretations (i.e., natural language concepts) of the subset of predictive neurons. To generate interpretations of a neuron, we prompt an LLM (GPT-4o in our experiments) with example texts from a range of neuron activations and instruct it to identify a natural language concept that is present in high-activation texts and absent in low-activation texts. We then generate multiple interpretations by rerunning this procedure several times at temperature 0.7, and choose the best interpretation according to a measure of fidelity we describe below.
We test a variety of approaches for generating interpretations. In our final procedure, we sample high- and low-activating texts from the neuron’s positive activation distribution, with the sampling bins given in Section B.2. (We restrict our sample to texts that activate the neuron to ensure some examples contain the concept.) Treating these as true positives and true negatives, respectively, we prompt an LLM to generate a natural language concept that distinguishes 10 samples of each class. We then define an interpretation’s fidelity as its F1 score in this prediction task, using an LLM (GPT-4o-mini in our experiments) to annotate 100 positive and 100 negative samples for the presence of the concept.
@@ 5.2 Metrics
  • Surface Similarity: For each matched pair, we prompt GPT-4o to assess whether the hypotheses are the same, related, or distinct, with corresponding scores of 1.0, 0.5, 0.0. To improve stability, we sample 5 outputs at temperature 0.7 and average them. We report the mean value of this score across the $H$ pairs. (Table 4 shows examples; Figure 5 provides the prompt.)
@@
The first and second metrics are taken directly from Zhong et al. (2024); they assess whether the inferred hypotheses qualitatively and quantitatively match the ground-truth. The third metric assesses overall prediction quality. Some hypotheses may not match a reference but still be predictive, so this metric rewards any interpretable hypotheses which contribute predictive value, even if they are not optimal.
@@
  • Breadth: We fit a multivariate logit (for binary labels) or OLS (continuous) regressing the target variable against all of the hypothesis annotations, and report the count of hypotheses which have a significant, nonzero coefficient [^4] We use a Bonferroni-corrected $p$-value threshold of 2.5e-3.. This metric benefits distinct hypotheses and penalizes redundancy. Breadth is important since the utility of hypotheses is often determined by criteria beyond predictiveness. For example, a hypothesis may be more useful if it is connected to past theories or suggests a new direction for investigation. More breadth increases the likelihood that a hypothesis satisfying these criteria is produced.
@@ 5.3 Baselines
BERTopic (Grootendorst, 2022) is a neural topic model which produces topics by clustering text embeddings. To produce $H$ hypotheses from BERTopic, texts are clustered into topics, and we fit an $L_{1}$-regularized model to predict $Y$ from topic ID; like our method, we set the penalty such that there are exactly $H$ nonzero coefficients. We assign each topic a natural language label by prompting an LLM with a sample of documents and top terms (prompt in Figure 7).
@@ 6.2 Real-World Datasets
Compared to baselines, HypotheSAEs produces the most hypotheses which are significant in a multivariate regression. Importantly, this evaluation reflects successful generalization: the hypotheses are learned on the training and validation sets, and they remain significantly associated with the target variable when GPT-4o-mini annotates them on the heldout set. Across the three datasets, 45 out of 60 candidate hypotheses are significant, substantially ahead of the 24, 23, and 20 for NLParam, BERTopic, and HypoGeniC respectively.
When evaluating overall predictive performance, HypotheSAEs beats baselines on 8 of 9 comparisons, demonstrating strongly predictive hypotheses. The exception is that HypoGeniC achieves 2% higher $R^{2}$ on Yelp. However, HypoGeniC’s most predictive hypotheses, such as expresses disappointment with the overall dining experience, essentially restate the rating prediction task. While LLMs can perform this prediction well, these hypotheses do not reveal specific insights for the task. In contrast, our hypotheses—in Table 6 for Yelp—tend to be more specific: we find that negative reviews mention food poisoning, rude and aggressive staff, or dishonest business practices.
@@ 6.4 Costs
Table 3: Runtimes and costs for hypothesis generation on Congress. We report millions of tokens for GPT-4o (feature labeling) and GPT-4o-mini (annotation) respectively as {# 4o} / {# 4o-mini}. Input and output token counts are summed (>>>99% are input). Sae-NoVal is our method without the label validation step (i.e., one label is generated and used per neuron). Costs are in USD and based on OpenAI pricing as of 01/2025.
@@ 7 Conclusion
We propose HypotheSAEs, a scalable method to generate interpretable hypotheses from black-box representations. On three synthetic and three real-world social science tasks, HypotheSAEs outperforms a topic model and state-of-the-art LLM baselines, at less than 10$\times$ the time and cost of the latter. The method applies naturally to many tasks in social science (Ziems et al., 2024), healthcare (Hsu et al., 2023; Robitschek et al., 2025; Pierson et al., 2025), and biology (Vig et al., 2021). More generally, there are many applications where we can accurately predict a target variable, but producing new scientific insights remains difficult. Our work helps bridge this gap.
@@
Code is available on GitHub, and can be installed via pip with pip install hypothesaes. A notebook to reproduce experimental results in the paper is available on the repository. All data used in the paper are available on HuggingFace. A demo to visualize all SAE neurons and how well they predict the target for Headlines, Yelp, and Congress is available at https://hypothesaes.org/.
@@ Acknowledgements
We thank Rishi Jha and Hal Triedman for helpful discussion about sparse autoencoders, Anna Lyubarskaja for helpful discussion about Proposition 3.1, Ruiqi Zhong for helpful discussion about NLParam, Neil Movva and OpenAI for LLM inference credits, Serina Chang for useful comments, Sammi Cheung for writing feedback, and members of the Pierson and Garg groups for comments and discussions.
@@ Appendix C Supporting Experiments
    Train an SAE on Yelp (Step 1), and identify the top-20 predictive neurons via Lasso (Step 2).
@@ Appendix D Hyperparameters for Baseline Methods
Appendix E Proofs

### v3

@@ 1 Introduction
Automated approaches to answering these questions have the potential to significantly expand and accelerate scientific discovery (Ludwig & Mullainathan, 2023). Indeed, multiple lines of work—spanning decades of research—have sought to extract text features that can predict a target variable (Blei et al., 2003; Monroe et al., 2009). Recent LLM-based hypothesis generation methods are especially exciting since they operate directly at the level of human-interpretable natural language concepts (Zhou et al., 2024; Batista & Ross, 2024). Concretely, hypothesis generation methods take a dataset of texts linked to a target variable (e.g., headlines and their engagement level) and hypothesize natural language concepts that predict the target variable (e.g., mentions being surprised or shocked or asks a question to the reader).
@@ 2.1 Hypothesis Generation
Recently, several LLM-based methods aim to automate interpretable concept discovery from text datasets. Grootendorst (2024) prompts LLMs to generate cluster labels, Pham et al. (2024) use LLMs for end-to-end topic modeling in natural language, and Lam et al. (2024) build an LLM-driven interface to discover topics guided by user input. Another literature uses LLMs to propose concepts in bottleneck models (Koh et al., 2020): Ludan et al. (2024) prompt LLMs with a task description to generate predictive concepts, and Feng et al. (2024) use LLMs to provide a prior distribution over concepts. Sun et al. (2024) and Zhong et al. (2024) also use LLMs to propose concepts, guided by neurons from a supervised model. Finally, Zhou et al. (2024) and Batista & Ross (2024) focus specifically on hypothesis generation: LLMs propose natural language descriptions of predictive relationships, which can then be tested for generalization.
@@ 3 Theoretical Framework
Intuitively, the result shows that if $\hat{Z}$ is a high-fidelity predictor of $Z$, then it will also have a similar separation score. A full proof is given in Appendix F. [^2] There are two main observations that enable us to show the result. First, we focus on $|\mathbb{E}[Y|\hat{Z}=1]-\mathbb{E}[Y|Z=1]|,$ and show that if the false negative rate (FNR) is greater than the false discovery rate (FDR), this value is bounded by the FNR; otherwise, it is bounded by the FDR. Second, we show roughly that as long as $\Pr[\hat{Z}=1],\Pr[Z=1]<\Pr[\hat{Z}=0],\Pr[Z=0],$ then $|\mathbb{E}[Y|\hat{Z}=0]-\mathbb{E}[Y|Z=0]|<|\mathbb{E}[Y|\hat{Z}=1]-\mathbb{E}[Y|Z=1]|.$
@@ 4 Methods
In the previous section, we established theoretically that we can generate hypotheses by learning interpretable neurons, identifying which neurons are predictive, and then generating high-fidelity interpretations of the neurons. We now describe our method, HypotheSAEs, which accomplishes these steps in practice.
Given a dataset of training examples $\{(x_{i},y_{i})\}_{i\in[N]},$ where $x_{i}$ is the input text and $y_{i}$ is the target variable annotation, HypotheSAEs outputs $H$ natural language concepts that serve as hypotheses. The goal is for these natural language concepts to predict the target variable.
@@
$z_{i}$ $=\operatorname{ReLU}(\operatorname{TopK}(W_{\operatorname{enc}}(e_{i}-b_{\mathrm{pre}})+b_{\operatorname{enc}})),$ (5)
@@
Empirically, the dimensions of $z_{i}$ are often highly interpretable, corresponding to human concepts. $M$ sets the total number of concepts learned across the entire dataset, and $k$ sets the number of concepts that can be used to reconstruct each instance. The output of this first step is an $N\times M$ activation matrix, $Z_{\text{SAE}}$.
@@
In this step, we generate high-fidelity interpretations (i.e., natural language concepts) of the subset of predictive neurons. To generate interpretations of a neuron, we prompt an LLM with example texts from a range of neuron activations and instruct it to identify a natural language concept that is present in high-activation texts and absent in low-activation texts. We generate multiple interpretations by running this procedure several times, and choose the best interpretation according to a measure of fidelity we describe below.
We test a variety of approaches for generating interpretations inspired by recent literature on automated interpretability (Bills et al., 2023; Templeton et al., 2024). In our final procedure, we sample high- and low-activating texts from the neuron’s positive activation distribution, with the sampling percentile bins given in Section B.2. (Compared to zero-activating texts, texts with a low but nonzero activation act as hard negatives, improving the generated concept’s precision.) Treating these as true positives and true negatives, respectively, we prompt an LLM to generate a natural language concept that distinguishes 10 samples of each class. We then define an interpretation’s fidelity as its F1 score in this prediction task, using an LLM to annotate 100 positive and 100 negative samples for the presence of the concept. We use GPT-4o with temperature 0.7 to generate concepts and GPT-4o-mini with temperature 0 for concept annotation.
@@ 5.2 Evaluation Metrics
  • Surface Similarity: For each matched pair, we prompt GPT-4o to assess whether the hypotheses are the same, related, or distinct, with corresponding scores of 1.0, 0.5, and 0.0. To improve stability, we sample 5 outputs at temperature 0.7 and average them. We report the mean value of this score across the $H$ pairs. (Table 4 shows examples; Figure 6 provides the prompt.)
@@
The first and second metrics are taken directly from Zhong et al. (2024); they assess whether the inferred hypotheses quantitatively and qualitatively match the ground-truth. The third metric assesses overall prediction quality. Some hypotheses may not match a reference but still be predictive, so this metric rewards any interpretable hypotheses that contribute predictive value, even if they are not optimal.
@@
  • Breadth: We fit a multivariate logit (for binary labels) or OLS (continuous) regressing the target variable against all of the hypothesis annotations, and report the count of hypotheses that have a significant, nonzero coefficient [^4] We use a Bonferroni-corrected $p$-value threshold of 2.5e-3.. This metric benefits distinct hypotheses and penalizes redundancy. Breadth is important since the utility of hypotheses is often determined by criteria beyond predictiveness. For example, a hypothesis may be more useful if it is connected to past theories or suggests a new direction for investigation. More breadth increases the likelihood that a hypothesis satisfying these criteria is produced.
@@
  • Qualitative Annotation: We recruit three computational science researchers to rate each hypothesis for (1) helpfulness and (2) interpretability (details in App. E). These attributes are inspired by prior work (Lam et al., 2024).
@@ 5.3 Baselines
BERTopic (Grootendorst, 2022) is a neural topic model that produces topics by clustering text embeddings. To produce $H$ hypotheses from BERTopic, texts are clustered into topics, and we fit an $L_{1}$-regularized model to predict $Y$ from topic ID; like our method, we set the penalty such that there are exactly $H$ nonzero coefficients. We assign each topic a natural language label by prompting an LLM with a sample of documents and top terms (prompt in Figure 8).
@@ 6.2 Real-World Datasets
Compared to baselines, HypotheSAEs produces the most hypotheses which are significant in a multivariate regression. Importantly, this evaluation reflects successful generalization: the hypotheses are learned on the training and validation sets, and they remain significantly associated with the target variable when annotated by GPT-4o-mini on the heldout set. Across the three datasets, 45 out of 60 candidate hypotheses are significant, substantially ahead of the 24, 23, and 20 for NLParam, BERTopic, and HypoGeniC respectively.
When evaluating overall predictive performance, HypotheSAEs beats baselines on 8 of 9 comparisons, demonstrating strongly predictive hypotheses. The exception is that HypoGeniC achieves 2% higher $R^{2}$ on Yelp. However, HypoGeniC’s most predictive hypotheses, such as expresses disappointment with the overall dining experience, essentially restate the rating prediction task. While most LLMs can perform such a prediction, these hypotheses do not reveal specific insights about the task. In contrast, our hypotheses—in Table 6 for Yelp—tend to be more specific: we find that negative reviews mention food poisoning, rude and aggressive staff, or dishonest business practices.
HypotheSAEs’s outputs are rated as qualitatively helpful and interpretable.
The results of our qualitative evaluation (Figure 2) confirm the above intuition: most of the predictive concepts generated by HypotheSAEs are rated helpful (24 out of 30) and interpretable (29 out of 30) [^5] To reduce annotator workload, our qualitative evaluation only included results from Headlines and Congress, since these are more well-studied social science tasks than Yelp.. Several hypotheses from NLParam and HypoGeniC, despite being predictive, do not pass these checks. The total count of hypotheses that are predictive, helpful, and interpretable generated by HypotheSAEs (23) far exceeds any baseline (BERTopic is second-best with 7).
Figure 2: HypotheSAEs produces a larger number (left) and percentage (right) of significant hypotheses that are rated helpful and interpretable. We report majority vote of 3 ratings per hypothesis.
6.3 Non-SAE Ablations
To isolate the value of the SAE, we conduct ablation experiments. Because our method differs from baselines in multiple ways—e.g., our neuron interpretation procedure differs from that of NLParam—it is worth asking what gain the SAE specifically provides. To answer this question, we run the following ablations on the real-world datasets:
  • Embedding: We skip training an SAE and directly select 20 features, via Lasso, from text embeddings. The rest of our method is unchanged.
  • PCA: Same as Embedding, but we first use PCA to reduce the embedding dimension from 1536 down to 256.
  • Bottleneck: Inspired by bottleneck models (Yang et al., 2023), we fit a neural network to predict $y$ using text embeddings as input and a single hidden layer with 20 neurons. We generate hypotheses by interpreting the hidden neurons (in the same way we interpret SAE neurons).
We show results in Table 7. In addition to the count of significant hypotheses and overall predictiveness, we also report average neuron fidelity.
We establish two benefits of the SAE: the neuron interpretations are (1) higher fidelity and (2) cover a wider variety of non-redundant concepts. Regarding (1), interpretations of SAE neurons achieve a mean F1 of 0.84, compared to 0.64 for the best ablation (bottleneck). Regarding (2), we see improvements to downstream metrics: the SAE produces many more significant hypotheses (45/60 across the three datasets, vs. 30/60 for the best ablation) and achieves slightly higher predictiveness. Qualitatively, the SAE neurons yield more specific interpretations, while the ablations tend to yield generic ones. Both benefits (1) and (2) are consistent with the SAE producing monosemantic neurons, which are easy to interpret and specific enough to avoid redundancy.
@@ 6.5 Costs
HypotheSAEs is much more efficient than NLParam and HypoGeniC. Despite the fact that these methods are trained on less data—20K examples vs. 114K for HypotheSAEsHypotheSAEs is $\sim$30-50$\times$ faster and 10-50$\times$ cheaper. This is because NLParam and HypoGeniC score hypotheses by annotating every example in the dataset. As a result, their LLM annotation requirements scale with both the number of hypotheses $H$ and dataset size $N$, while HypotheSAEs’s scales only with $H$.
@@
Table 3: Runtimes and costs for hypothesis generation on Congress. We report millions of tokens for GPT-4o (concept generation) and GPT-4o-mini (concept annotation) respectively as {# 4o} / {# 4o-mini}. Input and output token counts are summed (>>>99% are input tokens). Sae-NoVal is our method without the fidelity validation step (i.e., for each neuron, we generate only one label instead of choosing from three based on fidelity). Costs are in USD and based on OpenAI pricing as of 01/2025.
@@ 7 Conclusion
We propose HypotheSAEs, a scalable method to generate interpretable hypotheses from black-box text representations. On three synthetic and three real-world tasks, HypotheSAEs outperforms state-of-the-art topic modeling and LLM baselines, at less than 10$\times$ the time and cost of the latter. Beyond the tasks we study, the method applies naturally to the many settings in the social sciences where a large text corpus is labeled with a target variable of interest (Ziems et al., 2024; Bollen et al., 2011; Card et al., 2022; Breuer et al., 2025). In future work, an exciting direction is extending our method to explain non-text datasets, like images (Dunlap et al., 2024) and proteins (Vig et al., 2021), as well as text datasets from specialized domains like healthcare (Hsu et al., 2023; Robitschek et al., 2025). More generally, there are many applications where we can accurately predict a target variable, but producing new scientific insights remains difficult. Our work helps bridge this gap.
@@
[…]
@@ Acknowledgements
Thanks to Rishi Jha, Hal Triedman, Anna Lyubarskaja, Ruiqi Zhong, Serina Chang, and Sammi Cheung for helpful discussions. Thanks to Sidhika Balachandar, Divya Shanmugam, and Gabriel Agostini for qualitative annotation. Thanks to the ICML reviewers for helpful suggestions (specifically, to include SAE ablations and qualitative evaluation). Thanks to Neil Movva and OpenAI for LLM inference credits.
Funding: RM is supported by NSF DGE #2146752. NG is supported by NSF CAREER #2339427, and Cornell Tech Urban Tech Hub, Meta, and Amazon research awards. JK is supported by a Simons Collaboration grant and a grant from the MacArthur Foundation. EP is supported by a Google Research Scholar award, an AI2050 Early Career Fellowship, NSF CAREER #2142419, a CIFAR Azrieli Global scholarship, a gift to the LinkedIn-Cornell Bowers CIS Strategic Partnership, the Survival and Flourishing Fund, Open Philanthropy, and the Abby Joseph Cohen Faculty Fund.
@@ Appendix C Supporting Experiments
    Train an SAE (Step 1), and identify the top-20 predictive neurons via Lasso (Step 2).
@@ Appendix D Hyperparameters for Baseline Methods
Appendix E Qualitative Evaluation
To confirm that our hypotheses pass a “sniff test,” we recruit three computational social science researchers to perform a qualitative evaluation. These researchers are not authors on our study. We collect ratings for two attributes, following Lam et al. (2024): (1) helpfulness and (2) interpretability. We explain these to the raters as follows:
  1. 1.
    Helpful: Does this hypothesis help you understand what predicts the target variable? If you were studying this dataset, is this a hypothesis you would want to and could explore further? Rate 1 if yes, 0 if no or only a little.
  2. 2.
    Interpretable: When you read the hypothesis, is it clear what it means? Can you easily and unambiguously apply it to new examples? Rate 1 if yes, 0 if no or it would often be subjective.
We collect 3 annotations for each of the significant hypotheses outputted by any method on the Headlines and Congress datasets. (We removed the Yelp dataset to reduce annotator load.) Hypotheses are shown to annotators in random order, i.e., they were not grouped by method. We aggregate labels by taking the majority label.
Appendix F Proofs